OpenStudy (anonymous):

If you get paid $8.00 per hour at your job and insurance is $200 a month, how many hours will you have to work to pay for insurance?

OpenStudy (anonymous):

25 hours per month

OpenStudy (rsadhvika):

\(\large 200 \text{ dollars} \times \frac{1 hour}{8 \text{ dollars}}\)

OpenStudy (rsadhvika):

\(\large 200 \cancel{\text{ dollars}} \times \frac{1 hour}{8 \cancel{\text{ dollars}}}\)

OpenStudy (rsadhvika):

\(\large \cancel{200}^{25} \cancel{\text{ dollars}} \times \frac{1 hour}{\cancel{8} \cancel{\text{ dollars}}}\)

OpenStudy (rsadhvika):

\(\large 25 \text{ hours}\)

OpenStudy (marissalovescats):

25 hours.